OpenStudy (anonymous):

5 is what percent of 60

OpenStudy (anonymous):

To solve any problem like this just divide 5 by 60 so you get 0.08333333333

OpenStudy (anonymous):

can you not 60/5

OpenStudy (anonymous):

\[\Huge 8 \dfrac{1}{3} \%\]

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

since 1/4" = 10' on scale drawing what does 1/2" equal

OpenStudy (anonymous):

Multiply both sides by 2; 2(1/4") = 2(10') = 1/2" = ???

OpenStudy (anonymous):

can you help me to do it in proportions please

OpenStudy (anonymous):

can someone help me find the circumcenter for s(0,1) t(4,7) and U(8,-3)

OpenStudy (anonymous):

HELP!!!